OpenStudy (anonymous):

a box contain 9 novels, 7 biographies, and 5 reference books. Alyssa randomly chooses a book, then she randomly chooses another book without replacing the first book what is the probability that she chooses two biographies ? write as decimal

OpenStudy (kirbykirby):

P(choosing a biography) = 7/21, since there are 21 books in total and 7 of them are biographies. Once you selected a biography, there are only 6 left, and 20 books total, so the probability of choosing a second biography is 6/20 Hence, P(choosing biography one time AND choosing a biography a second time) = P(choosing biography one time) * P(choosing biography second time) = 7/21 * 6/20 = 1/10 = 0.1